# Algorithm Design and Prototype Validation for Reconfigurable Intelligent Sensing Surface: Forward-Only Transmission

### v2

@@ Abstract
Sensing-assisted communicatiion, DOA estimation, software-defined radio (SDR), over-the-air, prototyping.
@@ I Introduction
Notation: $\mathbf{I}_{M}$ denotes the $M\times M$ identity matrix, and $\mathbf{1}_{M}$ represents the $M\times 1$ column vector of all ones. The notation $[\cdot]_{i}$ refers to the $i$-th element of a vector, while $[\cdot]_{i,j}$ denotes the $(i,j)$-th element of a matrix. The imaginary unit is represented by $\mathbbm{i}=\sqrt{-1}$. The Euclidean norm is denoted by |||||$\cdot$||| | ⋅ | |, and the absolute value is denoted by ||$\cdot$|| ⋅ |. The function $\text{diag}(\cdot)$ constructs a diagonal matrix. The operators $(\cdot)^{T}$ and $(\cdot)^{H}$ denote the transpose and conjugate transpose, respectively. Finally, $\mathcal{CN}$ indicates the circularly symmetric complex Gaussian distribution.
@@ II-B Near-field Model
$d_{\text{Tx2R}}^{i,m,k}=\left|\left|\mathbf{b}_{\text{Tx},i,m}-\mathbf{b}_{\text{RISS},k}\right|\right|,$ (7)
$d_{\text{R2Rx}}^{k}=\left|\left|\mathbf{b}_{\text{RISS},k}-\mathbf{b}_{\text{Rx}}\right|\right|.$ (8)
@@
×$\qquad\qquad\qquad\qquad\qquad\times\left[\bar{\mathbf{h}}\right]_{k}\phi_{k}\left[\bar{\mathbf{G}}_{i}\right]_{k,m}\left[\bar{\mathbf{v}}_{i}\right]_{m}+n_{z}$
@@
Observe from Eq. (6) and Eq. (11) that the far-field and near-field models share similar expressions [^4] In practice, when the Tx and Rx are sufficiently far from the RIS, the waves transition from a near-field spherical wavefronts to far-field planar wavefronts, making the near-field and far-field models effectively equivalent. Generally, the far-field assumption holds when $R>\frac{2D^{2}}{\lambda}$. Here, $D$ represents the maximum dimension of the antenna, typically its diameter or maximum length. $R$ and $\lambda$ denote the distance between Tx/Rx and RIS and wavelength, respectively.. However, the specific channel expressions (i.e., Eq. (1)-(2) and Eq. (9)-(10)) differ. We will compare the disparities between the two channels through theoretical algorithm design and practical examination in the subsequent sections.
@@ III-A Far-field RISS-Aid Communications
$\mathbf{v}_{i}=\sqrt{P_{i}}\frac{\bm{\beta}^{\dagger}(\varpi_{G,i})}{||\bm{\beta}(\varpi_{G,i})||},$ (13)
where $P_{i}$ is the transmit power and |||||$\cdot$||| | ⋅ | | denotes the Euclidean norm.
@@
[…]
@@
$\bm{\mathcal{C}}_{i}=\mathbf{c}_{i}^{H}\mathbf{c}_{i},\forall i\in\{1,2\},$ (16d)
@@
[…]
@@
[…]
@@
where $\mathbf{V}_{t}$ denotes the matrix composed of eigenvectors corresponding to $[\lambda_{1},\ldots,\lambda_{N-1}]$, and $\lambda_{1}<\lambda_{2}<\ldots<\lambda_{N}$ represent the ascending eigenvalues of $\bm{\mathcal{C}}_{2}^{t}$ in the $t^{\text{th}}$ iteration. $r$ is the positive relaxation variable. As $r$ approaches zero, the undesired eigenvalue $[\lambda_{1},\cdots,\lambda_{N-1}]$ are forced to become zeros, which implies that $\bm{\mathcal{C}}_{2}^{t}$ satisfies the rank-one constraint. $\epsilon_{l}$ and $\varsigma$ represent gradually increasing auxiliary variables that control the rate at which $r$ decreases. Note that (P1.2) is a convex optimization problem, which can be solved by standard optimization method, such as the interior-point method. The IRM algorithm based solution is summarized in Algorithm 1.
@@
6:     $\epsilon_{t+1}=\epsilon_{t}^{\varsigma}$;
@@ III-B Near-field RISS-Aid Communications
×$\times\sum_{k\in N}\text{exp}\left(\frac{-\mathbbm{i}2\pi}{\lambda}\left(d_{\text{Tx2R}}^{i,m,k}+d_{\text{R2Rx}}^{k}\right)+\mathbbm{i}\varphi_{k}\right)+n_{z},$ (20)
@@
×$\quad\times\sum_{k\in N}\text{exp}\left(\frac{-\mathbbm{i}2\pi}{\lambda}\left(d_{\text{Tx2R}}^{2,k}+d_{\text{R2Rx}}^{k}\right)+\mathbbm{i}\varphi_{k}\right).$ (21)
Represent $y_{2}/\sqrt{\varrho_{\text{Tx2R},2}\varrho_{\text{R2Rx}}}$ with $S_{2}$, we have
$S_{2}=\sum_{k\in N}\text{exp}\left(\frac{-\mathbbm{i}2\pi}{\lambda}\left(d_{\text{Tx2R}}^{2,k}+d_{\text{R2Rx}}^{k}\right)+\mathbbm{i}\varphi_{k}\right).$ (22)
@@
$S_{1}=\sum_{k\in N}\text{exp}\left(\frac{-\mathbbm{i}2\pi}{\lambda}\left(d_{\text{Tx2R}}^{1,k}+d_{\text{R2Rx}}^{k}\right)+\mathbbm{i}\varphi_{k}\right).$ (23)
@@
[…]
s.t. $S_{i}=\sum_{k\in N}\text{exp}\left(\frac{-\mathbbm{i}2\pi}{\lambda}\left(d_{\text{Tx2R}}^{i,k}+d_{\text{R2Rx}}^{k}\right)+\mathbbm{i}\varphi_{k}\right)$
$\qquad\qquad\qquad\qquad\qquad\quad,\forall i\in\{1,2\},$ (24a)
@@
0:    Optimal passive phase shift matrix $\bm{\Theta}$.
@@
5:        Find the optimal $\phi_{k}^{*}$ that maximizes the objective function (P2) while keeping the other $\{\phi_{n}\}^{N}_{n=1,n\neq k}$ fixed;
6:        Update $\phi_{k}=\phi_{k}^{*}$;
@@ III-C Robust Design
where $\hat{\phi}{{}^{\text{azi}}_{G,1}}$ denotes the estimated incident angle of Tx. 1, and $\delta>0$ represents the boundary of the DOA estimation error. The fundamental robust design approach for both far-field and near-field scenarios involves dividing the interval $[\hat{\phi}{{}^{\text{azi}}_{G,1}}-\delta,\hat{\phi}{{}^{\text{azi}}_{G,1}}+\delta]$ into $L$ sub-intervals with a step size of $2\delta/{(L-1)}$, denoted by $\hat{\phi}{{}^{\S}_{1,l}},\forall l\in L$, and applying the suppression constraints to all angles within the interval.
@@ III-C1 Robust Design for Far-field Model
[…]
@@ III-C2 Robust Design for Near-field Model
s.t. $d_{\text{Tx2R},l}^{1,k}=\left|\left|\mathbf{b}_{\text{Tx},1,l}-\mathbf{b}_{\text{RISS},k}\right|\right|,$ (29a)
@@
where $S_{1,l}=\sum_{k\in N}\text{exp}\left(\frac{-\mathbbm{i}2\pi}{\lambda}\left(d_{\text{Tx2R},l}^{1,k}+d_{\text{R2Rx}}^{k}\right)+\mathbbm{i}\varphi_{k}\right)$. The detailed solution is outlined in Algorithm 2.
@@ III-D Computational Complexity and Convergence Analysis
The convergence behavior of both algorithms is illustrated in Fig. 3. It is observed that both algorithms converge within a finite number of iterations. Furthermore, the number of required iterations increases when the interference suppression constraints become more stringent, e.g., $\tau=-20$ dB and $\eta=5$. Therefore, appropriate interference suppression strategies can be selected based on the practical scenario to balance the trade-off between computational efficiency and system performance.
@@ IV-A Hardware Introductions
This section introduces the hardware employed within the system and delineates its implementation throughout. Fig. 4 presents an overview of the system’s hardware, while Table-II elaborates on the detailed parameters and functions of these hardware components. Note that we utilize an additional external clock (i.e., GPSDO) to synchronize local frequencies across different USRPs. Additional hardware parameters and descriptions, including the detailed specifications of the 2-bit RIS, the three-dimensional structural model, the $S_{11}$ reflection coefficient curve, and the radiation pattern of the sensing antenna array, are available in [39], which may offer further insights.
@@ IV-C1 DOA Estimation
where $\mathbf{E}_{s}\in\mathbb{C}^{N_{a}\times 2}$ and $\mathbf{E}_{z}\in\mathbb{C}^{N_{a}\times(N_{a}-2)}$ denote the signal and noise subspaces, respectively, spanned by their corresponding eigenvectors. $\bm{\Xi}_{s}$ and $\bm{\Xi}_{z}$ denote the corresponding eigenvalues. Due to the orthogonality between $\bm{\alpha}\left(\phi_{\text{Tx2R}}^{1}\right),\forall\in\{1,2\}$ and $\mathbf{E}_{z}$, the pseudo spectrum of MUSIC algorithm  [^8] The ROOT MUSIC [37] algorithm follows a process similar to that of the MUSIC algorithm, with the key distinction lying in the final step, where roots of the equation are sought instead of employing a spectral search as shown in Eq. (38). can be formulated as
@@ IV-C2 Signal Recognition
Figure 6: The structure of SRnet. The layer labeled ”Conv-k-s” signifies that this convolutional layer incorporates $k$ kernels, each of size $s$. All convolutions with a size of 3 are separable. The asterisk ()(*)( ∗ ) indicates that this layer employs a stride of size 2. Batch normalization (BN) is implemented subsequent to each convolution, followed by activation using rectified linear units (ReLUs).
@@ V Experiment Results
In this section, we practical verify the performance of RISS-assisted system. The entire system including two single horn-antenna Tx (i.e., $M=1$) and one single horn-antenna Rx, which are both controlled by the USRP X300 and transmit/receive carrier QPSK signals, as demonstrated in Section IV-A to IV-B. The location of the center of the RISS is set to $[0,0,0]$. The distance between Tx/Rx and the center of the RISS is set to $d_{c}=3.5$ meters, and the Tx/Rx and RISS are on the same plane, which means that $\phi^{\text{ele}}=\pi/2$, thus, the location of the Tx and Rx is set to $\mathbf{b}_{\text{Tx},i}=[d_{c}\sin(\phi^{\text{azi}}_{i}),0,d_{c}\cos(\phi^{\text{azi}}_{i})],\forall i\in\{1,2\}$ and $\mathbf{b}_{\text{Rx}}=[0,0,d_{c}]$. In the experiments, we can obtain $\phi^{\text{azi}}_{i}$ from the DOA algorithm, e.g., MUSIC and ROOTMUSIC, and the identification can be also obtained and determine which Tx is the target source, as depicts in Section IV-C. The interference suppression factors are set to $\tau=0.1$ (i.e., $-10$ dB) and $\eta=1$, respectively [^10] Note that excessively large values of $\tau$ and $\eta$ may lead to a greater emphasis on interference suppression, potentially at the expense of target signal enhancement. Therefore, when selecting the values of both parameters, it is essential to strike a balance between interference suppression and the enhancement of the target signal, in accordance with the specific requirements of the system.. Moreover, the Error Vector Magnitude (EVM) is adopted as a key metric to assess the modulation accuracy in digital communication systems. Typically expressed as a percentage, EVM quantifies the deviation between the received signal $s_{received}$ and the ideal reference signal $s_{ideal}$. It is mathematically defined as
@@
where RMS()($\cdot$)( ⋅ ) denotes the root mean square operation. The RMS values for the error vector and the ideal signal are given by
@@ V-B Sensing-assisted Communication for Signal Enhancement
Fig. 12b and Fig. 12d further illustrate that sensing errors have a slight impact on the communication performance of the target signal. This is because the main-lobe possesses an inherent width, as demonstrated in Fig. 2 and discussed in Section III-C. To further illustrate the phenomenon, the normalized beampattern measurements are presented in Fig. 13. Observed from that the main-lobe exhibits robustness for the target signal. Additionally, a side-lobe at $-20^{\circ}$ is evident in both the near-field and far-field models, which motivates the exploration of robust interference suppression schemes in the following section.

### v6

@@ Abstract
Sensing-assisted communication, DOA estimation, software-defined radio (SDR), over-the-air, prototyping.
@@ I Introduction
Notation: $\mathbf{I}_{M}$ denotes the $M\times M$ identity matrix, and $\mathbf{1}_{M}$ represents the $M\times 1$ column vector of all ones. The notation $\cdot]_{i}$ refers to the $i$-th element of a vector, while $\cdot]_{i,j}$ denotes the $(i,j)$-th element of a matrix. The imaginary unit is represented by $\mathbbm{i}=\sqrt{-1}$. The Euclidean norm is denoted by $||\cdot||$, and the absolute value is denoted by $|\cdot|$. The function $\text{diag}(\cdot)$ constructs a diagonal matrix. The operators $(\cdot)^{T}$ and $(\cdot)^{H}$ denote the transpose and conjugate transpose, respectively. Finally, $\mathcal{CN}$ indicates the circularly symmetric complex Gaussian distribution.
@@ II-B Near-field Model
[…]
@@
$\qquad\qquad\qquad\qquad\qquad\times\left[\bar{\mathbf{h}}\right]_{k}\phi_{k}\left[\bar{\mathbf{G}}_{i}\right]_{k,m}\left[\bar{\mathbf{v}}_{i}\right]_{m}+n_{z}$
@@
Observe from Eq. (6) and Eq. (11) that the far-field and near-field models share similar expressions [^4] In practice, when the Tx and Rx are sufficiently far from the RIS, the waves transition from a near-field spherical wavefronts to far-field planar wavefronts, making the near-field and far-field models effectively equivalent. Generally, the far-field assumption holds when $R>\frac{2D^{2}}{\lambda}$. Here, $D$ represents the maximum dimension of the antenna, typically its diameter or maximum length. $R$ and $\lambda$ denote the distance between Tx/Rx and RIS and wavelength, respectively.. However, the specific channel expressions (i.e., Eq. (1)-(2) and Eq. (9)-(10)) differ. We will compare the disparities between the two channels through theoretical algorithm design and practical examination in the subsequent sections [^5] While our current study focuses on realizing interference suppression and signal enhancement through a convex optimization framework, the Fourier optics approach [30] provides complementary insights by examining the RIS as a spatial filter through its transfer function..
@@ III-A Far-field RISS-Aid Communications
[…]
where $P_{i}$ is the transmit power and $||\cdot||$ denotes the Euclidean norm.
@@
$\text{(P1): }\max_{\bm{\Theta}}\quad$ $\text{trace}(\bm{\mathcal{C}}_{2}\bm{\mathcal{A}}_{2})$ (16)
@@
$\bm{\mathcal{C}}_{i}=\mathbf{c}_{i}^{H}\mathbf{c}_{i}\succeq\mathbf{0},\forall i\in\{1,2\},$ (16d)
@@
$\text{(P1.1): }\max_{\bm{\Theta}_{G}}\,\,\,$ $\text{trace}(\bm{\mathcal{C}}_{2}\bm{\mathcal{A}}_{2})$ (17)
@@
$\text{(P1.2): }\min_{\bm{\Theta}_{G}^{t}}\,\,\,$ $-\text{trace}(\bm{\mathcal{C}}_{2}^{t}\bm{\mathcal{A}}_{2})+\epsilon_{t}r$ (19)
@@
where $\mathbf{V}_{t}\in\mathbb{C}^{N\times N-1}$ denotes the matrix composed of eigenvectors corresponding to $\lambda_{1},\ldots,\lambda_{N-1}]$, and $\lambda_{1}<\lambda_{2}<\ldots<\lambda_{N}$ represent the ascending eigenvalues of $\bm{\mathcal{C}}_{2}^{t}$ in the $t^{\text{th}}$ iteration. $r$ is the positive relaxation variable. As $r$ approaches zero, the undesired eigenvalue $\lambda_{1},\cdots,\lambda_{N-1}]$ are forced to become zeros, which implies that $\bm{\mathcal{C}}_{2}^{t}$ satisfies the rank-one constraint. $\epsilon_{l}$ and $\varsigma$ represent gradually increasing auxiliary variables that control the rate at which $r$ decreases. Note that (P1.2) is a convex optimization problem, which can be solved by standard optimization method, such as the interior-point method. The IRM algorithm based solution is summarized in Algorithm 1.
@@
6:  $\epsilon_{t+1}=\epsilon_{t}\varsigma$;
@@ III-B Near-field RISS-Aid Communications
$\times\sum_{k\in N}\text{exp}\left(\frac{-\mathbbm{i}2\pi}{\lambda}\left(d_{\text{Tx2R}}^{i,m,k}+d_{\text{R2Rx}}^{k}\right)+\mathbbm{i}\varphi_{k}\right)+n_{z},$ (20)
@@
$\quad\times\sum_{k\in N}\text{exp}\left(\frac{-\mathbbm{i}2\pi}{\lambda}\left(d_{\text{Tx2R}}^{2,k}+d_{\text{R2Rx}}^{k}\right)+\mathbbm{i}\varphi_{k}\right).$ (21)
Represent $\left|y_{2}/\sqrt{\varrho_{\text{Tx2R},2}\varrho_{\text{R2Rx}}}\right|$ with $S_{2}$, we have
$S_{2}=\left|\sum_{k\in N}\text{exp}\left(\frac{-\mathbbm{i}2\pi}{\lambda}\left(d_{\text{Tx2R}}^{2,k}+d_{\text{R2Rx}}^{k}\right)+\mathbbm{i}\varphi_{k}\right)\right|.$ (22)
@@
$S_{1}=\left|\sum_{k\in N}\text{exp}\left(\frac{-\mathbbm{i}2\pi}{\lambda}\left(d_{\text{Tx2R}}^{1,k}+d_{\text{R2Rx}}^{k}\right)+\mathbbm{i}\varphi_{k}\right)\right|.$ (23)
@@
(P2): $\max_{\varphi_{k},k\in N}\,S_{2}-\eta S_{1}$ (24)
s.t. $S_{i}=\left|\sum_{k\in N}\text{exp}\left(\frac{-\mathbbm{i}2\pi}{\lambda}\left(d_{\text{Tx2R}}^{i,k}+d_{\text{R2Rx}}^{k}\right)+\mathbbm{i}\varphi_{k}\right)\right|$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad,\forall i\in\{1,2\},$ (24a)
@@
0:   Optimal passive phase shift matrix $\varphi_{k},k\in N$.
@@
5:   Find the optimal $\varphi_{k}^{*}$ that maximizes the objective function (P2) while keeping the other $\{\varphi_{n}\}^{N}_{n=1,n\neq k}$ fixed;
6:   Update $\varphi_{k}=\varphi_{k}^{*}$;
@@ III-C Robust Design
where $\hat{\phi}{{}^{\text{azi}}_{G,1}}$ denotes the estimated incident angle of Tx. 1, and $\delta>0$ represents the boundary of the DOA estimation error. The fundamental robust design approach for both far-field and near-field scenarios involves dividing the interval $\hat{\phi}{{}^{\text{azi}}_{G,1}}-\delta,\hat{\phi}{{}^{\text{azi}}_{G,1}}+\delta]$ into $L$ sub-intervals with a step size of $2\delta/{(L-1)}$, denoted by $\hat{\phi}{{}^{\S}_{1,l}},\forall l\in L$, and applying the suppression constraints to all angles within the interval.
@@ III-C1 Robust Design for Far-field Model
$\text{(P1.3): }\min_{\bm{\Theta}_{G}^{t}}\,\,\,$ $-\text{trace}(\bm{\mathcal{C}}_{2}^{t}\bm{\mathcal{A}}_{2})+\epsilon_{t}r$ (28)
@@ III-C2 Robust Design for Near-field Model
where $S_{1,l}=\left|\sum_{k\in N}\text{exp}\left(\frac{-\mathbbm{i}2\pi}{\lambda}\left(d_{\text{Tx2R},l}^{1,k}+d_{\text{R2Rx}}^{k}\right)+\mathbbm{i}\varphi_{k}\right)\right|$. The detailed solution is outlined in Algorithm 2.
@@ III-D Computational Complexity and Convergence Analysis
The detailed convergence analyses are available in Appendix A, and the convergence behavior of both algorithms is illustrated in Fig. 3. Observed that both algorithms converge within a finite number of iterations. Furthermore, the number of required iterations increases when the interference suppression constraints become more stringent, e.g., $\tau=-20$ dB and $\eta=5$. Therefore, appropriate interference suppression strategies can be selected based on the practical scenario to balance the trade-off between computational efficiency and system performance.
@@ IV-A Hardware Introductions
This section introduces the hardware employed within the system and delineates its implementation throughout. Fig. 4 presents an overview of the system’s hardware, while Table-II elaborates on the detailed parameters and functions of these hardware components. Note that we utilize an additional external clock (i.e., GPSDO) to synchronize local frequencies across different USRPs. Additional hardware parameters and descriptions, including the detailed specifications of the 2-bit RIS, the three-dimensional structural model, the $S_{11}$ reflection coefficient curve, and the radiation pattern of the sensing antenna array, are available in Appendix B, which may offer further insights.
@@ IV-C1 DOA Estimation
where $\mathbf{E}_{s}\in\mathbb{C}^{N_{a}\times 2}$ and $\mathbf{E}_{z}\in\mathbb{C}^{N_{a}\times(N_{a}-2)}$ denote the signal and noise subspaces, respectively, spanned by their corresponding eigenvectors. $\bm{\Xi}_{s}$ and $\bm{\Xi}_{z}$ denote the corresponding eigenvalues. Due to the orthogonality between $\bm{\alpha}\left(\phi_{\text{Tx2R}}^{i}\right),\forall i\in\{1,2\}$ and $\mathbf{E}_{z}$, the pseudo spectrum of MUSIC algorithm  [^9] The ROOT MUSIC [38] algorithm follows a process similar to that of the MUSIC algorithm, with the key distinction lying in the final step, where roots of the equation are sought instead of employing a spectral search as shown in Eq. (38). can be formulated as
@@ IV-C2 Signal Recognition
Figure 6: The structure of SRnet. The layer labeled ”Conv-k-s” signifies that this convolutional layer incorporates $k$ kernels, each of size $s$. All convolutions with a size of 3 are separable. The asterisk ()(*) indicates that this layer employs a stride of size 2. Batch normalization (BN) is implemented subsequent to each convolution, followed by activation using rectified linear units (ReLUs).
@@ V Experiment Results
In this section, we practical verify the performance of RISS-assisted system. The entire system including two single horn-antenna Tx (i.e., $M=1$) and one single horn-antenna Rx, which are both controlled by the USRP X300 and transmit/receive carrier QPSK signals, as demonstrated in Section IV-A to IV-B. The location of the center of the RISS is set to $[0,0,0]$. The distance between Tx/Rx and the center of the RISS is set to $d_{c}=3.5$ meters, and the Tx/Rx and RISS are on the same plane, which means that $\phi^{\text{ele}}=\pi/2$, thus, the location of the Tx and Rx is set to $\mathbf{b}_{\text{Tx},i}=[d_{c}\sin(\phi^{\text{azi}}_{i}),0,d_{c}\cos(\phi^{\text{azi}}_{i})],\forall i\in\{1,2\}$ and $\mathbf{b}_{\text{Rx}}=[0,0,d_{c}]$. In the experiments, we can obtain $\phi^{\text{azi}}_{i}$ from the DOA algorithm, e.g., MUSIC and ROOTMUSIC, and the identification can be also obtained and determine which Tx is the target source, as depicts in Section IV-C. The interference suppression factors are set to $\tau=0.1$ (i.e., 10-10 dB) and $\eta=1$, respectively [^11] Note that excessively large values of $\tau$ and $\eta$ may lead to a greater emphasis on interference suppression, potentially at the expense of target signal enhancement. Therefore, when selecting the values of both parameters, it is essential to strike a balance between interference suppression and the enhancement of the target signal, in accordance with the specific requirements of the system.. Moreover, the Error Vector Magnitude (EVM) is adopted as a key metric to assess the modulation accuracy in digital communication systems. Typically expressed as a percentage, EVM quantifies the deviation between the received signal $s_{received}$ and the ideal reference signal $s_{ideal}$. It is mathematically defined as
@@
where RMS()($\cdot$) denotes the root mean square operation. The RMS values for the error vector and the ideal signal are given by
@@ V-B Sensing-assisted Communication for Signal Enhancement
Fig. 12b and Fig. 12d further illustrate that sensing errors have a slight impact on the communication performance of the target signal. This is because the main-lobe possesses an inherent width, as demonstrated in Fig. 2 and discussed in Section III-C. To further illustrate the phenomenon, the normalized beampattern measurements are presented in Fig. 13. Observed from that the main-lobe exhibits robustness for the target signal. Additionally, a side-lobe at $20^{\circ}$ is evident in both the near-field and far-field models, which motivates the exploration of robust interference suppression schemes in the following section.